Please show your work and explain for me so I can understand better please. 1. Complete the following Balance Sheet Equations: Assets Liability Stockholders Equity a. 245000 130000 ? b. 318000 Liabili
Answer #1

The Balance sheet equation is

Assets = Liabilities + Stockholder's Equity

now lets solve one by one

a)Assets = Liabilities + Stockholder's Equity

giving value to the equation, we will get

245000 = 130000+ Stockholder's Equity

So Stockholder's Equity = 245000-130000 = 115000

b)In this problem consider Stockholder's Equity as "X"

so Stockholder's Equity = X

Liabilities is twice of Stockholder's Equity

So Liabilities = 2X

giving value to the equation, we will get

318000 = 2X + X

So 3X = 318000

X = 106000

So Stockholder's Equity = 106000

Liabilities = 2X = 212000

c) Liabiities is half of Stockholder's Equity

Stockholder's Equity = 168000

So Liabilities = 168000/2 = 84000

Assets = Liabilities + Stockholder's Equity

giving value to the equation, we will get

Assets = 84000+ 168000 = 252000

Please free to comment for further clarification ,and please upvote if it was helpful